Hello Expert Sir pease help me and send the answer of the following answer.
where sine and cosine curve meet
1
Expert's answer
2011-12-01T08:16:35-0500
Let's solve equation sinx = cosx. As we can see, the equation is linear, in sin x and cos x. The way of solving this kind of equation is to divide entire equation with "cos x".

(sin x / cos x) - 1 = 0

Instead of the ratio "sin x / cos x', we'll use the resulting function, tg x.

The obtained equation is an elementary equation.

tg x - 1 = 0

tg x = 1

x= arctg 1 +k*pi

So, x= pi/4 + k*pi.